OpenStudy (anonymous):

Given sin B = .65 find the angle B in radians. Round your answer to the nearest hundredth. please help me with this question

OpenStudy (souvik):

\[B=\sin^{-1} (.65)\]

OpenStudy (anonymous):

would it be 1.67 radians ?

OpenStudy (souvik):

what is B in degree?

OpenStudy (anonymous):

doesnt show it just tells me to solve for it and the multiple questions are A=1.67 B=1.34 C=.98 D=.71

OpenStudy (souvik):

B= 40.54 degree in radians it is 0.71... do u know what is the relation between degree and radians?

OpenStudy (whpalmer4):

Here's a hint: 

OpenStudy (whpalmer4):

pi/6 = 0.524 radians pi/4 = 0.785 radians pi/3 = 1.0472 radians which unit circle value of the three has sin(x) closest to 0.65?

OpenStudy (whpalmer4):

sin(pi/6) = 1/2 = 0.5 sin(pi/4) = sqrt(2)/2 = 0.707 sin(pi/3) = sqrt(3)/2 = 0.866

OpenStudy (whpalmer4):

From this you should be able to see that the answer choice closest to pi/4 is the one to choose.

OpenStudy (anonymous):

thank you

OpenStudy (whpalmer4):

of course with a calculator you could just take the arcsin of 0.65 and get the answer directly, but where's the challenge in that? (well, making sure your calculator is set in radian mode, I suppose)